OpenStudy (anonymous):

Please! Im terrible at breaking apart division with place value! 846 divided by 2 = ( + 40 + ) divided by 2 = (800+40+6) ( 2 + 2+2) = (400 + + 3) =

OpenStudy (turingtest):

I don't know if I understand you but 846=(800+40+6) 846/2=(800/2+40/2+6/2)=400+20+3=423